OpenStudy (anonymous):

a spring requires a force of 25 n to produce an extension of 2 cm in it. What is the work done in futher extending it b

OpenStudy (anonymous):

u know about integrals?

OpenStudy (anonymous):

easy way... k=25/2 if b is in cms......... w = 0.5*k*b*b total work = 0.5*k*(b+2)^2

OpenStudy (asnaseer):

rayonia: Welcome to OpenStudy! Since this is your first question I assume you are not familiar with the site yet. The question you have posted is more suited to the Physics group. Please Close this Question and then click on the blue "Find More Subjects" link at the top of this web page and join the Physics group. If you post this question there then you will have more chance of finding someone that is able to help. If you still find no one to help you (as some groups are quieter than others) then please feel free to post a link to your question in any of the chat windows and ask the users there for assistance.